OpenStudy (mollykenney):

I NEED HELP!!! convert the rectangular equation to a polar equation that expresses r in terms of theta x^2=16y r=?

OpenStudy (astrophysics):

Note, \[x=rcos \theta \] \[y = rsin \theta \] then we have \[r=\sqrt{x^2+y^2}\]

OpenStudy (zale101):

@mollykenney does it make sense? :)

OpenStudy (mollykenney):

so we have sqrt of x^2 + (-16)y = r?

OpenStudy (zale101):

The question wants you to convert this x^2=16y to polar equations. @Astrophysics already gave you what you need to know.

OpenStudy (zale101):

x^2=16y into polar coordinates and then from there you solve for r. When you convert it to polar coordinates, you will get r and theta as your new variables. Solve for r.

OpenStudy (zale101):

what does x equal to and y equal to in polar coordinates?

OpenStudy (mollykenney):

i don't know @Zale101 can we take it step by step I'm really confused with this concept

OpenStudy (astrophysics):

Would you agree, if \[x=rcos \theta \] then \[x^2 = r^2\cos^2 \theta \]

OpenStudy (astrophysics):

What would \[16y = ?\]

OpenStudy (reemii):

quick and dirty: \(x=r\cos\theta, y=r\sin\theta\), then \(16x^2 = y \iff 16 r^2\cos^2\theta = r\sin\theta\iff r(16r\cos^2\theta - \sin\theta)=0\) \(\iff r=0 \text{ or } r=\frac{\sin\theta}{16\cos^2\theta} \) So it's the relation \(r = \frac{\sin\theta}{16\cos^2\theta} \), \(\theta\in[0,\pi]\setminus\{\pi/2\}\).